Assume (1) a company’s plantwide predetermined overhead rate is $13.00 per direct labor-hour, and (2) its job cost sheet for Job X shows that this job used 18 direct labor-hours and incurred direct materials and direct labor charges of $500 and $360, respectively. If Job X contains 46 units, then its unit product cost is closest to:

  • $23.78.

  • $19.44.

  • $20.18.

  • $15.96.
Answer #1

Total applied overhead=Predetermined overhead rate*Actual direct labor hours

=18*13=$234

Add:Direct materials=$500

Add:Direct labor=$360

Total cost =$1094

Hence unit product cost=Total cost/units of job

=1094/46

=$23.78(Approx)